# Authentic Discrete Diffusion Model

### v1

@@ 4.3 Multi-Token Generation Task (Text Generation)
PDD-Generated captions 0.18
ADD-Generated captions 0.25
@@
We further investigate the potential of ADD in multi-token generation tasks, with a particular focus on image captioning as a representative text generation benchmark. Unlike single-label classification, captioning requires the model to generate coherent sequences of tokens conditioned on visual inputs, thereby testing its ability to model discrete sequential dependencies. This task is especially challenging for diffusion-based approaches, as it requires capturing both the syntactic structure of natural language and the semantic alignment with images.
A major difficulty in this setting is the relatively limited scale of available text supervision in datasets such as COCO, which provides only short captions per image. In contrast, state-of-the-art large language models are typically trained on massive corpora with billions of tokens using extensive computational resources. Consequently, training a competitive captioning model directly on COCO represents a stringent test of efficiency and adaptability. To address the scarcity of text tokens, we utilized CLIP Scores as the evaluation metric. This metric employs CLIP to extract image and text features, and calculates the similarity between these two features, providing an well-suited evaluation of semantic relevance. In addition, we present qualitative examples to illustrate the fluency and semantic accuracy of generated captions.
@@
For comparison, we benchmark our ADD against the masked diffusion framework, which we refer to as the “pseudo” discrete diffusion (PDD) model. PDD generates tokens by predicting masked entries in a partially observed sequence, but unlike ADD, it does not operate in a fully discrete diffusion process and therefore lacks consistent alignment with categorical token spaces.
@@
To evaluate the semantic relevance between the captions and the corresponding images quantitatively, we introduced a pre-trained CLIP model in the experimental section to extract features from both text and images and compute their similarity in Table 3. We evaluated the CLIP Scores for several combinations: genuine image-text pairs, mismatched pairs (where the text is grammatically correct but largely unrelated to the image), PDD-generated text paired with images, and ADD-generated text paired with images. The results clearly demonstrate that the captions predicted by ADD exhibit strong semantic alignment with the images.
@@
Beyond numerical evaluation, qualitative comparisons are provided in Fig. 7. Captions generated by ADD exhibit grammatical correctness, semantic coherence, and close alignment with both the input image content and human-annotated ground-truth captions. In contrast, outputs from PDD frequently suffer from broken syntax and poor semantic fidelity, often producing phrases that are disconnected from the visual scene. These observations reinforce our claim that ADD, by operating directly in the one-hot label space with timestep-conditioned cross-entropy–based training, preserves both the discrete structure of text tokens and their semantic alignment with conditioning inputs.
@@
Taken together, these results highlight two key insights. First, authentic discrete diffusion substantially improves text generation quality over pseudo-discrete alternatives. Second, our findings suggest that ADD can serve as a foundation for broader applications in multi-token generation, bridging the gap between discrete modeling in vision tasks and natural language generation. This dual capability further strengthens the generality of ADD as a unified framework for discrete generative modeling.
@@ 5 Conclusion
We presented Authentic Discrete Diffusion (ADD), a new framework that extends diffusion modeling to categorical domains while preserving its core generative principles. Unlike prior pseudo-discrete approaches that rely on embeddings or masking-based approximations, ADD operates directly in the one-hot space through a combination of cross-entropy alignment loss, $\arg\max$-based discretization, and iterative denoising. Our experiments on classification and text generation demonstrate that ADD achieves strong performance with minimal sampling steps, significantly outperforming pseudo-discrete baselines in both quantitative metrics (e.g., accuracy and perplexity) and qualitative evaluations. Ablation studies further highlight the indispensability of each component, confirming that cross-entropy alignment loss and classifier-free guidance are critical to making discrete diffusion both stable and effective. Overall, ADD establishes a principled and efficient pathway for applying diffusion models to symbolic data. We believe this work provides a foundation for future research on discrete generative modeling, with potential extensions to large-scale language tasks, multimodal reasoning, and structured prediction problems.
6 Future Work
While Authentic Discrete Diffusion (ADD) establishes a principled framework for discrete-space generative modeling, several promising directions remain open for future exploration.
Theoretical Foundations.
A natural extension of this work is to formalize the theoretical underpinnings of ADD. Current results demonstrate empirical stability and effectiveness, yet deeper analysis is needed to rigorously characterize its convergence behavior, sample complexity, and generalization bounds. Establishing guarantees on the approximation quality of $\arg\max$-based denoising, as well as the statistical efficiency of noise-schedule–aware cross-entropy weighting, would provide a stronger foundation for understanding ADD as a discrete analog of continuous diffusion. Moreover, connecting ADD to information-theoretic principles—such as entropy reduction in categorical spaces—could yield general laws governing discrete diffusion processes.
Scaling Laws.
Future applications of ADD span language, multimodal, and embodied domains. In large language models, ADD offers a discrete diffusion alternative to autoregressive training, with the potential to mitigate exposure bias, improve long-sequence stability, and enable efficient multi-token prediction. Extending ADD to multimodal LLMs could unify discrete-space denoising across text, vision, and structured categorical data, fostering categorical consistency and interpretability in cross-modal reasoning tasks such as grounding and retrieval. Beyond language and vision, ADD also provides a pathway toward large physical world models, enabling scalable and interpretable systems that integrate reasoning, perception, and action for embodied AI.

### v2

@@ 4.3 Multi-Token Generation (Text Generation)
MDD-Generated captions 0.18
TDD-Generated captions 0.25
@@
We further investigate the potential of TDD in multi-token generation tasks by evaluating on ImageNet-1K-VL-Enriched and MS-COCO captioning datasets, with a particular focus on image captioning as a representative text generation benchmark. Unlike single-label classification, captioning requires the model to generate coherent sequences of tokens conditioned on visual inputs, thereby testing its ability to model discrete sequential dependencies. This task is especially challenging for diffusion-based approaches, as it requires capturing both the syntactic structure of natural language and the semantic alignment with images.
We first present qualitative results on ImageNet-1K-VL-Enriched imagenet-1k-vl-enriched, which provides diverse image–text pairs for testing multi-token coherence. Representative captions on the validation split are shown in Fig. 8.
Figure 8: Qualitative comparison of image captioning between ground-truth annotations and TDD-generated captions. TDD demonstrates strong semantic understanding across diverse visual scenes, accurately identifying objects and scenes.
To mitigate the impact brought by the scarcity of text tokens, we utilized CLIP Scores as the evaluation metric. This metric employs CLIP to extract image and text features, and calculates the similarity between these two features, providing an well-suited evaluation of semantic relevance. In addition, we present qualitative examples to illustrate the fluency and semantic accuracy of generated texts.
@@
For comparison, we benchmark our TDD against the masked diffusion framework, which we refer to as the masked discrete diffusion (MDD) model. MDD generates tokens by predicting masked entries in a partially observed sequence, but unlike TDD, it does not operate in a fully discrete diffusion process and therefore lacks consistent alignment with categorical token spaces.
@@
To quantitatively assess the semantic relevance between captions and images, we employed a pre-trained CLIP model to extract features from both modalities and compute their similarity (Tab. 3). We evaluated CLIP Scores for genuine image-text pairs, mismatched pairs, MDD-generated captions, and TDD-generated captions. The results clearly show that TDD-generated captions exhibit strong semantic alignment with the corresponding images.
@@
Beyond numerical evaluation, qualitative comparisons in Fig. 7 show that TDD-generated captions are grammatically correct, semantically coherent, and well-aligned with both the input image and human-annotated ground-truth captions. In contrast, MDD outputs often exhibit broken syntax and poor semantic fidelity, with phrases disconnected from the visual content. These results support our claim that TDD, by operating directly in the one-hot label space, maintains both the discrete structure of text tokens and their semantic alignment with conditioning inputs.
@@
Taken together, these results highlight two key insights. First, TDD substantially improves text generation quality over masked-discrete alternatives. Second, our findings suggest that TDD can serve as a foundation for broader applications in multi-token generation, bridging the gap between discrete modeling in vision tasks and natural language generation. This dual capability further strengthens the generality of TDD as a unified framework for discrete generative modeling.
@@ 5 Conclusion
We introduced Tweedie Discrete Diffusion (TDD), a diffusion framework for categorical data that leverages cross-entropy alignment, $\operatorname*{arg\,max}$ discretization, and iterative denoising. Unlike masked-discrete methods, TDD denoises directly in one-hot space. Experiments demonstrate that TDD delivers strong performance with very few sampling steps, surpassing masked-discrete baselines. Ablations further show that cross-entropy alignment and classifier-free guidance are essential for stable discrete diffusion. Overall, TDD provides a principled route for applying diffusion models to symbolic domains, with promising extensions to large-scale language modeling.
Appendix A Implementation Details
A.1 Setup
Datasets and Augmentation.
As mentioned in the experimental section, our classification experiments are conducted on the ImageNet(5206848) dataset, incorporating a diverse set of advanced data augmentation strategies.
For superior classification performance, we employ a comprehensive data augmentation pipeline. During training, images are randomly resized and cropped to the target resolution, followed by horizontal flipping. We use the AutoAugment policy rand-m9-mstd0.5-inc1, which introduces a diverse set of photometric and geometric transformations. Color jitter is disabled in our default setting.
We further adopt Random Erasing with probability 0.25, using the pixel mode and a single erase block. All images are normalized using the standard ImageNet mean and standard deviation.
For regularization, we apply label smoothing with a smoothing factor of 0.1. In our primary experiments, Mixup and CutMix are disabled (mixup = 0, cutmix = 0). When enabled, mixing is performed with probability 1.0, switching to CutMix with probability 0.5, under a batch-level mixing strategy.
During evaluation, the input is resized using bicubic interpolation with a crop ratio, followed by a center crop to the target size. Finally, images are converted to tensors and normalized using the same ImageNet statistics as in training.
For text generation, TDD is trained on the ImageNet-1K-VL-Enriched(imagenet-1k-vl-enriched) dataset together or the COCO(lin2014microsoft) dataset.
For large-scale image–text pre-training in the captioning task, we adopt the ImageNet-1K-VL-Enriched dataset, which is built on the standard ImageNet-1K classification benchmark and further enriched with automatically generated captions, object-level bounding boxes and auxiliary metadata. Each sample comprises the original image, its ImageNet class label, a textual caption generated by BLIP2, and additional information that is not utilized in our experiments. This enriched corpus thereby transforms the original classification dataset into a versatile vision–language resource while retaining the taxonomy and visual diversity of ImageNet.
To complement ImageNet-1K-VL-Enriched with richly annotated everyday scenes and enable a more comprehensive and unbiased evaluation of TDD, we employ the COCO 2017 dataset. COCO (Common Objects in Context) consists of approximately 118 287 training images and 5 000 validation images. Each image is annotated with five human-written captions and object annotations across 80 common categories. In our experiments, we use the official caption annotations for training and validation. Owing to its emphasis on natural, multi-object and context-rich scenes, COCO 2017 provides an ideal benchmark for evaluating the semantic fidelity and linguistic diversity of our model.
Training Protocol.
TDD is trained from scratch for 500 epochs distributed across $8\times$ NVIDIA A100 80GB GPUs. All experiments are conducted using PyTorch 2.7.1 with CUDA 12.8.
Evaluation Metrics.
For ImageNet classification, we report Top-1 accuracy on the validation set.
For COCO captioning, our primary metric is CLIP Scores computed between image and text features; higher is better.
A.2 Samping Design
During sampling, we designed two distinct approaches for processing the model’s output at each iteration, as detailed below.
The first is argmax sampling, which directly selects the class with the highest probability from the output and projects it into a one-hot space. This can be viewed as a hard probability space.
The second approach is multinomial sampling, which first transforms the model’s output into a soft probability space via softmax, then employs the torch.multinomial function for probability-based sampling, and finally maps the result to a cleaner one-hot space.
A.3 Classifier-Free Guidance and Exponential Moving Average
Classifier-Free Guidance (CFG) and Exponential Moving Average (EMA) are two essential components in our generation pipeline. CFG enables controllable conditioning strength during sampling, while EMA stabilizes model parameters to improve sample quality and evaluation robustness. We describe both mechanisms and detail how they are implemented in our framework.
Classifier-Free Guidance.
CFG is a widely adopted technique in conditional generative modeling. Given a conditional prediction $x_{\text{cond}}$ and an unconditional prediction $x_{\text{uncond}}$, the final guided prediction is computed as:
$x_{\text{guided}}=x_{\text{uncond}}+\text{cfg}\cdot(x_{\text{cond}}-x_{\text{uncond}}),$ (1)
where $\text{cfg}\geq 1$ controls the guidance strength. A value of $\text{cfg}=1$ corresponds to unguided sampling, while larger values amplify semantic alignment with conditioning information.
In our implementation, CFG is applied at every autoregressive step during token sampling. The model learns a dedicated unconditional class embedding, allowing conditional and unconditional forward passes to be computed jointly.
Input: Model $f$, initial latent $x$, guidance scale cfg
for $t=1$ to $T$ do
   Compute $u_{t}=f(x_{t},\text{null})$;
   Compute $c_{t}=f(x_{t},\text{cond})$;
 $\gamma_{t}\leftarrow s(t)$;
 $x_{t}=u_{t}+\gamma_{t}(c_{t}-u_{t})$;
   Sample token and update $x_{t+1}$;
Algorithm 1 Sampling with CFG
This sampling design enables TDD to balance semantic fidelity and sample diversity while adapting guidance intensity throughout the iterative process.
Exponential Moving Average.
EMA maintains a smoothed copy of model parameters, which is updated after every training step:
$\theta_{\text{EMA}}\leftarrow\alpha\,\theta_{\text{EMA}}+(1-\alpha)\,\theta,$ (2)
where $\alpha$ is typically close to 1 (we use $\alpha=0.9999$). This moving average reduces the stochastic noise of parameter updates and produces more stable convergence.
In practice, our system keeps two parameter sets: the raw parameters $\theta$ and the smoothed EMA parameters $\theta_{\text{EMA}}$. During evaluation, we switch to the EMA parameters, as they consistently yield better captioning quality and reduced variance across runs.
Input: Model parameters $\theta$, EMA parameters $\theta_{\text{EMA}}$, decay rate $\alpha$
for each training iteration do
   Compute gradients and update $\theta$;
 for each parameter $\theta_{i}$ do
    $\theta_{\text{EMA},i}\leftarrow\alpha\cdot\theta_{\text{EMA},i}+(1-\alpha)\cdot\theta_{i}$;
return $\theta_{\text{EMA}}$
Algorithm 2 EMA Update During Training
EMA therefore serves two purposes: (1) mitigating overfitting to noisy updates during training, and (2) ensuring that sampling and evaluation rely on a more stable, lower-variance parameter trajectory.
Appendix B Additional Experiments
B.1 Extended Analysis of TDD’s Reverse Denoising Dynamics
Figure 9: Additional qualitative examples for TDD. For each input image (left), we show TDD’s textual predictions at multiple diffusion timesteps (right), ranging from highly corrupted discrete token states at large timesteps to coherent natural-language descriptions at $t=0$. The visualization reveals how semantic structure emerges gradually throughout the reverse denoising trajectory.
In this appendix, we provide an extended qualitative analysis of the reverse denoising behavior exhibited by our TDD (Tweedie Discrete Diffusion) framework. The examples in Figure 9 show how TDD transforms heavily corrupted discrete token states into coherent natural-language captions as the diffusion process proceeds toward lower timesteps. These visualizations expose the full semantic refinement path taken by the model, offering interpretability insights not observable from final-step predictions alone.
At large timesteps (e.g., $t=90$), the model operates under high levels of corruption, producing token sequences with extremely high entropy. Because TDD injects Gaussian noise into a relaxed embedding representation before projecting back to the discrete one-hot space, the decoded outputs contain fragmented or fabricated lexical elements, exhibiting no meaningful alignment with the visual input. These states reflect the structural priors of the tokenizer rather than any grounded semantics.
As the reverse diffusion trajectory advances to intermediate timesteps (around t=20t=201515), semantically relevant anchors begin to emerge. These anchors typically manifest as nouns corresponding to visually salient objects, but remain intermixed with substantial residual noise. This stage reveals the earliest point at which TDD begins to integrate visual cues into the discrete denoising pathway. Notably, object-centric tokens tend to surface before relational or structural linguistic elements, reflecting their greater robustness to noise.
In the later stages of the process ($t=10\rightarrow 0$), TDD consolidates the partially formed semantic structure into coherent sentences. Once key object terms stabilize, the model progressively recovers syntactic elements, attribute modifiers, and connective words, ultimately producing clean and grammatically valid captions. The discrete projection step—enforced at every iteration—encourages tokens to remain stable once they have emerged, reducing semantic blending and ensuring sharp, unambiguous outputs at $t=0$.
The examples also highlight several characteristic failure modes that complement the quantitative findings in the main paper. TDD sometimes captures the dominant object while overlooking secondary entities (e.g., detecting a person but not a skateboard), suggesting that finer-grained cues may be underrepresented in the visual encoder. Attribute-level distinctions—such as differentiating “boy” from “man”—may also be lost under moderate noise, leading to subtle semantic drift during intermediate timesteps. Additionally, temporary deviations into irrelevant token clusters occasionally appear mid-trajectory but typically correct themselves as denoising proceeds.
Overall, these qualitative observations emphasize the interpretability advantages of discrete diffusion. TDD provides direct access to intermediate linguistic states, enabling detailed inspection of how semantic information accumulates and stabilizes across timesteps. The analysis here demonstrates a consistent progression from high-entropy noise to structured language and identifies promising directions for improving fine-grained semantic fidelity in future iterations of the model.
B.2 Token-Level Denoising Behavior in TDD
Figure 10: Token-level trajectories during TDD denoising. Each heatmap corresponds to a single token from the earlier qualitative examples and visualizes the categorical score assigned to several competing token classes across the diffusion sampling steps. These trajectories reveal how individual token semantics transition from high-entropy noise at large timesteps to stable linguistic interpretations as the model approaches $t=0$.
In Figure 9, we illustrated the sentence-level denoising behavior of our TDD framework. To further analyze how individual tokens evolve during the reverse diffusion process, we provide in Figure 10 a set of heatmaps that visualize the categorical score trajectories of selected tokens across the full sampling path. Each heatmap corresponds to a single token taken from the earlier qualitative examples and plots the model’s predicted scores for several candidate token categories as a function of the diffusion timestep. These trajectories reveal fine-grained semantic dynamics that are not captured by sentence-level visualizations alone.
At large timesteps, the model operates under heavy corruption, producing token-level predictions with high entropy. For some tokens, the model assigns nearly all probability mass to a dominant but incorrect category, showing that early reverse diffusion steps are largely governed by noise and embedding priors rather than grounded semantics. As denoising proceeds, the probability gradually shifts toward semantically meaningful categories, revealing the incremental emergence of linguistic structure.
Across the three heatmaps, we observe distinct patterns of token behavior. The first heatmap illustrates a token undergoing clear semantic correction: its dominant category transitions from an incorrect class at early timesteps toward the correct class as visual and contextual evidence accumulates. This behavior corresponds to tokens in the sentence-level visualizations that transform from incoherent fragments into meaningful words. The second heatmap depicts a token whose prediction remains stable throughout the entire trajectory. Such tokens are typically function words or high-confidence content words that are easily recoverable even under heavy noise, resulting in minimal fluctuation. The third heatmap shows a token experiencing strong temporal competition between multiple candidate categories. The model alternates among several plausible semantic hypotheses during intermediate steps before eventually converging to a final class. This reflects tokens in the sentence-level examples that exhibit ambiguity—often involving fine-grained distinctions such as age, object type, or subtle attributes—before stabilizing near the end of the sampling process.
These token-level trajectories demonstrate the interpretability advantages of discrete diffusion. Because TDD performs explicit denoising in one-hot token space, the intermediate states remain directly interpretable as categorical distributions, enabling us to observe how semantics are formed, refined, and occasionally revised as diffusion progresses. This analysis complements the sentence-level qualitative results and provides a deeper understanding of how discrete diffusion reconstructs linguistic structure from noise, shedding light on both the strengths of TDD’s semantic refinement process and its characteristic failure modes, such as late-stage ambiguity resolution or transient class switching.